In this talk we will define arithmetic hyperbolic 3-manifolds, a class of 3-manifolds built from number fields. We will discuss how geometric questions about arithmetic manifolds, such as the volume or the lengths of shortest geodesics, can be answered using algebraic tools and invariants.